# Efficient Discriminative Joint Encoders for Large Scale Vision-Language Reranking

### v1

@@ 5.2 Main results
We first examine whether EDJE, when considered as a lightweight joint encoder with minimal capacity, can substantially improve retrieval performance over embedding-based pipelines. To this end, we deploy the local variant as a top-$k$ reranker: for each embedding model tested, EDJE reuses its vision backbone and pairs it with MiniLM as the shared language encoder. We evaluate zero-shot retrieval performance on Flickr30k with standard text-to-image and image-to-text retrieval tasks. The results are summarized in Table˜1.
EDJE boosts retrieval performance across all tested embedding models, emphasizing the potential of integrating rerankers to existing retrieval pipelines. Specifically, we observe massive gains for the original CLIP Radford et al. (2021) model, with Recall@1 improvements of up to 15% for image retrieval and 10% for text retrieval. Noticeable gains are also obtained for the SigLIP2 backbone, despite it being a highly optimized state-of-the-art embedding model. The improvements for DFN and MetaCLIP are less noticeable; however, DFN relies on a filtering network fine-tuned on Flickr.
We next assess EDJE when considered as a practical alternative to prior joint encoders Li et al. (2021; 2022; 2023), fixing the visual backbone to SigLIP 2 with a resolution of $384^{2}$ to match their setup. To ensure fairness, we evaluate both local and token-compressed variants under a cached-vision regime. Namely, we assume that the visual features are precomputed, so that only the online joint-encoding part is considered allowing us to compare against previous methods. Under this setup we compare methods along several axes: retrieval accuracy (Recall@1 on Flickr, zero-shot, and COCO, fine-tuned, for both image-to-text and text-to-image), per-image storage (kilobytes), joint-encoder parameter count, online inference time (milliseconds for a batch of 64 on an A6000 GPU), and the amount of training data used. The results are summarized in Table˜2.
EDJE achieves a favorable accuracy–efficiency profile relative to existing joint encoders. The local variant matches prior work on Flickr (zero-shot) and remains competitive on COCO (fine-tuned), while using a much smaller joint-encoder (tens of millions of parameters rather than hundreds) and substantially lower online latency. Crucially, these gains come with affordable storage costs: even in its uncompressed form it may be suitable for some use-cases, and the token-compressed variant has minimal storage costs while preserving most of the retrieval accuracy.
@@ 5.3 Ablation Studies
We first analyze how varying the number of compressed tokens affects retrieval performance. Specifically, we evaluate Flickr30k zero-shot image retrieval using $\{32,64,128,256\}$ target tokens in the token-compression adapter (Figure˜4). As expected, increasing the number of tokens yields better performance, with a clear gap between the heavily compressed $32$-token variant and the uncompressed "local" variant ($576$ tokens). Notably, $64$ tokens strike an attractive balance between efficiency and retrieval quality.
@@
Finally, we ablate the pre-training objectives introduced in Section˜4. For the local variant, we compare: (i) optimizing image–text matching (ITM) alone, (ii) ITM combined with masked language modeling (MLM), and (iii) the full objective that further adds text-embedding recovery. Each auxiliary loss contributes positively, with the full objective delivering the strongest results. In addition, we evaluate the impact of local-to-compressed knowledge distillation, which provides further gains for compressed variants by transferring discriminative capacity from the local model. We refer the reader to Appendix˜B for the full results.
@@ 6 Discussion
We studied how to make joint vision–language rerankers practical at scale. We recognize visual feature extraction as the key bottleneck in existing joint encoders. We tackle this problem by introducing EDJE. The approach of EDJE is to precompute the vision tokens and compress them with a lightweight adapter in an offline manner, in addition to a compact joint encoder that can deliver high-throughput inference while retaining high performance.
Figure 5: Retrieval performance vs. reranking pool size. Robustness of local and 64-token variants under different pool sizes on Flickr.
@@
We think of this paper as a proof of concept that may inspire follow-up work; for instance, we did not cover multilingual-multimodal retrieval, which has drawn attention recently Thapliyal et al. (2022) or other modalities such as audio or video. More broadly, we believe joint encoders are largely underexplored; putting effort into improving them can benefit a variety of applications including zero-shot classification and filtering large paired datasets.
@@ Appendix A Training Hyperparameters
Table A summarizes the main hyperparameters used throughout pre-training and fine-tuning. We fix the language model to MiniLM-L12-H384-uncased and freeze the vision encoder in all runs. Unless otherwise stated, all experiments use in-batch negative mining with three negatives per sample, and a reranking pool size of $k=10$. These settings were chosen to balance training efficiency and retrieval quality and remained consistent across all backbones.
@@
Reranking pool size $k$ 10
@@
Appendix B Additional Ablation Studies
@@
Table 5 focuses on token compression and highlights the effect of applying distillation. Here, knowledge distillation provides additional improvements when compressing visual tokens.

### v2

@@ 5.2 Main results
We first examine whether EDJE, when considered as a lightweight joint encoder with minimal capacity, can substantially improve retrieval performance over embedding-based pipelines. To this end, we deploy the local variant as a top-$k$ re-ranker: for each embedding model tested, EDJE reuses its vision backbone and pairs it with MiniLM as the shared language encoder. We evaluate zero-shot retrieval performance on Flickr30k with standard text-to-image and image-to-text retrieval tasks. The results are summarized in Table˜1.
EDJE boosts retrieval performance across all tested embedding models, emphasizing the potential of integrating re-rankers to existing retrieval pipelines. Specifically, we observe massive gains for the original CLIP (Radford et al., 2021) model, with Recall@1 improvements of up to 15% for image retrieval and 10% for text retrieval. Noticeable gains are also obtained for the SigLIP2 backbone, despite it being a highly optimized state-of-the-art embedding model. The improvements for DFN and MetaCLIP are less noticeable; however, DFN relies on a filtering network fine-tuned on Flickr.
We next assess EDJE when considered as a practical alternative to prior joint encoders (Li et al., 2021; 2022; 2023), fixing the visual backbone to SigLIP 2 with a resolution of $384^{2}$ to match their setup. To ensure fairness, we evaluate both local and token-compressed variants under a cached-vision regime. Namely, we assume that the visual features are precomputed, so that only the online joint-encoding part is considered allowing us to compare against previous methods. Under this setup we compare methods along several axes: retrieval accuracy (Recall@1 on Flickr, zero-shot, and COCO, fine-tuned, for both image-to-text and text-to-image), per-image storage (kilobytes), joint-encoder parameter count, online inference time (milliseconds for a batch of 64 on an A6000 GPU), and the amount of training data used. The results are summarized in Table˜2.
EDJE achieves a favorable accuracy–efficiency profile relative to existing joint encoders. The local variant matches prior work on Flickr (zero-shot) and remains competitive on COCO (fine-tuned), while using a much smaller joint-encoder (tens of millions of parameters rather than hundreds) and substantially lower online latency. Crucially, these gains come with affordable storage costs: even in its uncompressed form it may be suitable for some use-cases, and the token-compressed variant has minimal storage costs while preserving most of the retrieval accuracy. Interestingly, we find out that quantizing the compressed tokens before storing them, then de-quantizing upon inference yields minimal performance degradation and can further improve storage-performance tradeoff. We refer the reader to Appendix˜H for more details.
@@ 5.3 Ablation Studies
Figure 5: Retrieval performance vs. re-ranking pool size. Robustness of local and 64-token variants under different pool sizes on Flickr.
@@
We begin by analyzing how varying the number of compressed tokens affects retrieval performance. Specifically, we evaluate Flickr30k zero-shot image retrieval using $\{32,64,128,256\}$ target tokens in the token-compression adapter (Figure˜5). As expected, increasing the number of tokens yields better performance, with a clear gap between the heavily compressed $32$-token variant and the uncompressed "local" variant ($576$ tokens). Notably, $64$ tokens strike an attractive balance between efficiency and retrieval quality. We additionally compare against simpler token-reduction strategies, providing alternative baselines to the proposed adapter; see Appendix˜G for details.
@@
Finally, we ablate the pre-training objectives introduced in Section˜4. For the local variant, we compare: (i) optimizing image–text matching (ITM) alone, (ii) ITM combined with masked language modeling (MLM), and (iii) the full objective that further adds text-embedding recovery. Each auxiliary loss contributes positively, with the full objective delivering the strongest results. We further evaluate the impact of local-to-compressed knowledge distillation, which provides further gains for compressed variants by transferring discriminative capacity from the local model. We also investigate cross-model negative selection to understand how EDJE behaves under different vision-encoder geometries. We refer the reader to Appendix˜B and Appendix˜E for more details.
5.4 Interpretable Semantics of Compressed Vision Tokens
To better understand the information preserved by the compressed visual tokens produced by EDJE, we aim to analyze their semantic structure regardless of any caption that may be equipped with the original image. We achieve this by inspecting the projection of each visual token into the language-model embedding space and assigning its nearest textual token from the LM vocabulary. Collecting these nearest neighbors (one per compressed tokens) reveals which language tokens the model most frequently associates with visual tokens, as depicted in Figure˜6.
Figure 6: Semantic structure of the 64-token compressed representation. Left: example image from the Flickr-30k test set. Right: nearest-neighbor textual tokens assigned to compressed vision tokens; word size reflects frequency in the token distribution.
The compressed tokens map to meaningful object and scene descriptors such as boulders, caves, glare, or trio, indicating that the adapter preserves important semantic information. In contrast, we find that interestingly many of the uncompressed 576-token representation map to a meaningless special tokens (unused80), suggesting that a large portion of native ViT tokens carry redundant content for retrieval. We refer the reader to Appendix˜D for further explanations and experiments.
@@ 6 Discussion
We studied how to make joint vision–language re-rankers practical at scale. We recognize visual feature extraction as the key bottleneck in existing joint encoders. We tackle this problem by introducing EDJE. The approach of EDJE is to precompute the vision tokens and compress them with a lightweight adapter in an offline manner, in addition to a compact joint encoder that can deliver high-throughput inference while retaining high performance.
@@
We think of this paper as a proof of concept that may inspire follow-up work; for instance, we did not cover multilingual-multimodal retrieval, which has drawn attention recently (Thapliyal et al., 2022) or other modalities such as audio or video. More broadly, we believe joint encoders are largely underexplored; putting effort into improving them can benefit a variety of applications including zero-shot classification and filtering large paired datasets.
@@ Appendix A Training Hyperparameters
Table A summarizes the main hyperparameters used throughout pre-training and fine-tuning. We fix the language model to MiniLM-L12-H384-uncased and freeze the vision encoder in all runs. Unless otherwise stated, all experiments use in-batch negative mining with three negatives per sample, and a re-ranking pool size of $k=10$. These settings were chosen to balance training efficiency and retrieval quality and remained consistent across all backbones.
@@
Re-ranking pool size $k$ 10
@@
Appendix B Ablation study: training objectives
@@
Table 5 focuses on token compression and highlights the effect of applying distillation. Here, knowledge distillation provides additional improvements when compressing visual tokens. This section analyzes the interpretability properties of the compressed visual tokens produced by EDJE. Although the token compression module discards a large portion of the original vision encoder sequence, we find that the resulting compact representation retains coherent semantic structure. We study two aspects: (1) the emergence of high-level concepts within individual compressed tokens, and (2) the alignment between compressed visual tokens and text features used during retrieval.
@@
Appendix C Using EDJE in practice
We provide a pseudo-code for using EDJE in practical retrieval pipelines, combining a clip-like embedding model, a vector store and EDJE as an efficient re-ranking model. A pseudo-code for the indexing stage is given in Algorithm˜1 while the retrieval stage is summarized in Algorithm˜2.
Algorithm 1 Indexing (offline)
1:Images (images), vector store (store), CLIP-like vision encoder (vision_encoder), finetuned EDJE token adapter (adapter)
2:image_loader = DataLoader(images)
3:for image_batch in image_loader do
4:  // Extract both the usual embeddings and patch features in one pass
5:  features, embeddings = vision_encoder(image_batch)
6:  // Apply the token adapter on the encoders’ features
7:  features = adapter(features)
8:  // Store embeddings on RAM and features on disk
9:  store.insert(
10:   embeddings=embeddings,
11:   extra_data={"features": features}
12: )
13:end for
Algorithm 2 Retrieval
1:Query text (query), vector store (store), CLIP-like text encoder (text_encoder), finetuned EDJE re-ranker (re-ranker), re-ranking pool size (k)
2:re-ranked retrieval results
3:// Compute the usual query text-embedding
4:text_embedding = text_encoder(query)
5:// Retrieve candidates from vector store
6:candidates = store.knn_query(text_embedding, k=k)
7:// Compute re-ranking scores with EDJE
8:features = torch.cat([candidate["features"] for candidate in candidates], dim=0)
9:scores = re-ranker(query, features)
10:results = candidates[scores.argsort(descending=True)] return results
C.1 Token-Fetch I/O Considerations
Beyond compute, a practical deployment of EDJE must also account for the cost of fetching precomputed vision tokens from storage. In our experiments, each image is represented by 64 compressed tokens in BF16, corresponding to roughly $49$ kB per image. For a re-ranking pool of 5050k candidates, this amounts to reading approximately $2.46$ GB of data per query.
To estimate the expected I/O overhead, we benchmark two storage layouts on a PCIe 4.0 NVMe SSD: (i) a single contiguous NumPy array storing all image representations, and (ii) a more realistic memory-mapped array accessed via random indices. In the contiguous case, loading the full $2.46$ GB block takes around $0.39\pm 0.003$ s over 10 runs (approximately $6.3$ GB/s), which is in line with the advertised bandwidth of the SSD. With random access over a 100100k-image memory-mapped index (50k random entries), the same amount of data is loaded in $0.59\pm 0.04$ s.
These measurements indicate that, on modern local SSDs, the I/O cost of fetching compressed tokens is on the same order as the compute cost of the joint encoder and still allows end-to-end processing of tens of thousands of pairs per second. Note however that networked storage can exhibit substantially lower throughput, so we do not recommend it for EDJE deployments.
Appendix D Visual tokens interpretability
D.1 emerging vision-token semantics
We next study the semantic content of the visual tokens produced by EDJE in a caption-independent manner. The goal is to qualitatively understand what information the compressed visual tokens carry, regardless of any ground-truth caption that may or may-not be provided.
Given an image, let $\{v_{i}\}_{i=1}^{n}$ denote either the compressed visual tokens (e.g, 64 tokens from the token-compression adapter) or the full ViT sequence (e,g 576 tokens, transformed locally). We first project each $v_{i}$ into the language-model embedding space using the same projection as in the joint encoder. For every visual token $v_{i}$, we then retrieve its nearest language token $w_{i}$ from the LM vocabulary (using cosine similarity). Collecting these nearest neighbors $\{w_{i}\}_{i=1}^{n}$ over the Flickr30k test set allows us to analyze which language-tokens the model most frequently associates with visual tokens.
This analysis has two desirable properties: (i) it evaluates the semantics of visual tokens purely through the geometry of the joint embedding space, without using the paired captions for that image, and (ii) it yields interpretable text tokens that can be visualized as through frequency histograms. We perform this analysis for two concrete example images for the compressed tokens, as illustrated in Figure˜7.
Figure 7: Emerging token semantics for the 64-token compressed representation. Left (per row): example images from Flickr test set. Right (per row): textual tokens word cloud, size indicates frequency according to vision token nearest-neighbors histogram.
The compressed tokens exhibit stable and highly interpretable semantics. Tokens frequently correspond to concrete object categories (e.g., dog, collar, boulders, trio) and scene attributes (e.g., glare, shadows). Despite the large reduction in sequence length, the model preserves a rich vocabulary of visual cues. We than repeated the analysis using the "local" EDJE variant, as illustrated in Figure˜8.
Figure 8: Token semantics for the 576-local EDJE representation. Left (per row): example images from Flickr test set. Right (per row): textual tokens word cloud, size indicates frequency according to vision token nearest-neighbors histogram.
In this case, however, a large fraction of tokens map to a special vocabulary item, unused80, or to scattered low-frequency words, resulting in much less concentrated distributions. This suggests that many of the original ViT tokens have low semantic content in the joint space and are largely redundant for downstream retrieval.
D.2 Image-Caption Token Alignment
We complement the caption-independent analysis above with a quantitative study of how well the compressed visual tokens align with caption tokens. The goal is to measure to what extent cross-modal alignment is preserved when reducing the number of visual tokens from the full ViT sequence to the compressed EDJE representation.
For each image-caption pair, we first preprocess the caption by lowercasing all words and removing punctuation and standard English stopwords (e.g., "A child playing in the ocean." $\rightarrow$ "child playing ocean").
Let $\{t_{j}\}_{j=1}^{m}$ denote the remaining textual tokens and $\{v_{i}\}_{i=1}^{n}$ the set of visual tokens for the corresponding image, where $\{v_{i}\}$ is either the full (locally transformed) ViT sequence (e.g., $n=576$) or the compressed set produced by EDJE (e.g., $n=64$).
We embed both text tokens and visual tokens into the joint embedding space using the same projections as in the joint encoder, and compute a ColBERT-like alignment score [Khattab and Zaharia, 2020, Faysse et al., 2025]. For each textual token $t_{j}$, we compute its maximum cosine similarity to any visual token,
$s_{j}=\max_{1\leq i\leq n}\cos\bigl(f(t_{j}),g(v_{i})\bigr),$
where $f(\cdot)$ and $g(\cdot)$ are the corresponding text and vision projections. We then define the alignment score for the image–caption pair as the average of these maxima,
$S=\frac{1}{m}\sum_{j=1}^{m}s_{j}.$
Intuitively, $S$ measures how well each caption word can “find” at least one strongly related visual token in the image. Note that we didn’t directly optimize $S$ as a scoring criterion.
We compute the distribution of alignment scores $S$ over all Flickr30k test pairs for both the full 576-token representation and the 64-token compressed representation as depicted in Figure˜9.
Figure 9: Distribution of alignment scores for full token sequences and their token compressed counterparts. The histogram shows that compression preserves the overall alignment structure while slightly shifting the distribution toward lower but still tightly concentrated scores. This supports the claim that the compressed representations retain most of the semantic signal needed for retrieval.
The token-compressed representation achieves an average alignment score of $0.2516\pm 0.0491$, while the full-sequence representation achieves $0.2404\pm 0.0405$. The corresponding histograms overlap and exhibit very similar shapes, with the compressed representation showing a slightly higher mean. A two-sided paired t-test on the per-pair scores indicates that this difference is statistically significant ($p=3.0\times 10^{-153}$), with a small-to-moderate effect size ($d=0.39$). Overall, these results show that EDJE ’s 64-token representation preserves, and slightly improves, image-caption alignment relative to the full 576-token ViT sequence, despite using $9\times$ fewer visual tokens.
Appendix E Cross Model Negative Mining
In the main experiments, hard negatives for EDJE are mined using the same embedding model family as the underlying retriever (e.g., SigLIP2 for a SigLIP2-based EDJE). To assess the impact of this design choice, we additionally study cross-model negative mining, where the miner and the backbone come from different embedding models.
Concretely, we consider two settings: (i) using SigLIP2 as a negative miner for a CLIP-based EDJE, and (ii) using CLIP as a negative miner for a SigLIP2-based EDJE. In both cases, the joint encoder and token-compression components remain fixed; only the model used to select hard negatives during fine-tuning is changed.
Table 6: Using SigLIP2 as a negative miner for a CLIP based EDJE. Mining with SigLIP2 does not substantially affect CLIP performance.
Model Flickr Zero-Shot COCO Finetuned
Text-to-Image Image-to-Text Text-to-Image Image-to-Text
CLIP ViT L/14@336 81.9 93.8 54.6 70.7
+SigLIP2 L miner 81.8 93.5 54.6 70.8
Table 7: Using CLIP as a negative miner for a SigLIP2 based EDJE. Mining with CLIP significantly degrades SigLIP2 performance.
Model Flickr Zero-Shot COCO Finetuned
Text-to-Image Image-to-Text Text-to-Image Image-to-Text
SigLIP2 ViT L/16@384 87.8 96.5 64.9 81.0
+CLIP L miner 81.5 93.0 58.9 74.48
The results in Tables 6 and 7 reveal an interesting pattern. When SigLIP2 is used as a negative miner for a CLIP-based EDJE, performance changes only marginally: Flickr30k zero-shot R@1 remains essentially unchanged, and COCO fine-tuned scores are within a similar range. This suggests that in this setting, overall performance is largely limited by the vision backbone rather than by the precise geometry of the negative miner. In contrast, when CLIP is used as a miner for a SigLIP2-based EDJE, performance drops substantially on both Flickr30k and COCO, for both text-to-image and image-to-text retrieval. This degradation emphasizes the importance of selecting sufficiently hard negatives. Overall, these experiments support the practical guideline that hard-negative mining should be performed with a model that is at least as strong as, and well aligned with, the underlying vision-encoder used in EDJE.
Appendix F Full Dataset Retrieval Evaluation
To provide a more comprehensive evaluation, we adopt the full-dataset retrieval protocol used in LightningDOT [Sun et al., 2021], where retrieval is performed against all images and captions in the dataset, including the train/validation splits. This setting is considerably more challenging and better reflects real-world retrieval scenarios. We follow LightningDOT’s setup for both Flickr Full and COCO Full, and scale the re-ranking pool size to $100$ in all experiments.
Table 8: Flickr Full retrieval results under the LightningDOT setup. Retrieval is performed against all dataset’s images/captions as retrieved instances. EDJE is evaluated in a zero-shot setting and substantially outperforms LightningDOT in both text-to-image and image-to-text retrieval.
Model Text-to-Image Image-to-Text
R@5 R@10 R@20 R@5 R@10 R@20
LightningDOT 60.1 69.5 78.3 75.1 83.9 90.5
EDJE 78.32 84.54 89.58 92.4 95.9 97.7
Table 9: COCO Full retrieval results under the LightningDOT setup. Retrieval is performed against all dataset’s images/captions as retrieved instances. EDJE significantly outperforms LightningDOT across all recall levels and directions.
Model Text-to-Image Image-to-Text
R@5 R@10 R@20 R@5 R@10 R@20
LightningDOT 37.3 46.8 56.4 48.0 59.0 68.9
EDJE 52.23 60.55 68.08 69.86 76.96 82.64
As shown in Tables 8 and 9, EDJE substantially improves full-dataset retrieval performance over LightningDOT on both Flickr Full and COCO Full. On Flickr Full (zero-shot), EDJE yields large gains in both directions and at all recall levels (e.g., text-to-image R@5 improves from $60.1$ to $78.32$, and image-to-text R@5 from $75.1$ to $92.40$). On COCO Full (fine-tuned), EDJE again surpasses LightningDOT by a wide margin (e.g., text-to-image R@5 from $37.3$ to $52.23$, image-to-text R@5 from $48.0$ to $69.86$). These results confirm that EDJE remains effective in more realistic large-candidate retrieval scenarios.
Appendix G Token Compression Baselines
In this section we compare the proposed token-compression adapter in EDJE to several alternative ways of reducing the number of visual tokens produced by the SigLIP2 ViT-L encoder. All methods start from the same 576-token ViT sequence ($\times384$ resolution) and compress it to 64 tokens per image. We evaluate on Flickr30k (zero-shot) and COCO (fine-tuned).
We consider the following token-compression strategies:
  1. 1.
    Striding. A simple subsampling baseline that keeps every $9$th token from the $576$-token ViT sequence, yielding $64$ tokens in total $(576/9=64)$.
  2. 2.
    Token clustering. We run k-means++ over the $576$ visual tokens for each image to obtain $64$ clusters, and use the cluster centroids as compressed tokens.
  3. 3.
    Attention pruning. We compute the attention scores of each token in the last ViT layer and keep the $64$ most attended tokens (where attendance is averaged across heads).
For fairness, each configuration is pretrained and fine-tuned end-to-end with the same protocol as our token-compression model. Particularly, we distill from the uncompressed (576-token) teacher as in the main EDJE experiments.
Table 10: Comparison of token-compression strategies applied to the SigLIP2 ViT-L image encoder. We report Recall@1 on Flickr30k (zero-shot) and COCO (fine-tuned). SigLIP2 (Baseline) denotes the original embedding-only model. All other methods compress the 576 ViT tokens to 64 tokens per image and use the same joint encoder architecture.
Model Flickr Zero-Shot COCO Finetuned
Text-to-Image Image-to-Text Text-to-Image Image-to-Text
SigLIP2 (Baseline) 82.3 94.8
Striding 83.24 94.1 60.09 77.2
Clustering 85.66 96.1 63.31 79.76
Attention Pruning 82.4 93.7 58.6 76.4
EDJE 86.9 96.4 64.6 80.9
We observe that EDJE’s token-compression adapter consistently outperforms striding, clustering, and attention pruning across both datasets and directions (text-to-image and image-to-text). This emphasizes the superiority of the learned query-based adapter over generic pooling or pruning schemes.
Appendix H Quantizing the Vision Tokens
In all main experiments, the precomputed vision tokens are stored in BF16 (or FP16) format. Because EDJE is designed for large-scale retrieval with potentially billions of stored image representations, it is interesting to examine how far the storage footprint can be reduced via quantizing the numerical precision of the vision tokens.
We perform post-training quantization only on the precomputed vision tokens, while keeping the joint encoder itself in BF16. For each quantization type, the cached tokens are quantized on disk and de-quantized immediately before being fed to the joint encoder at inference time, with no additional finetuning.
We consider two representation families: (i) the full 576-token ViT sequence, and (ii) the 64-token compressed representation. For each family we evaluate three numeric formats: (a) BF16 (no quantization), (b) FP8-E4M3, and (c) FP4-E2M1. We measure Flickr30k zero-shot image retrieval performance (R@1) and compute the average storage per image in kilobytes, taking into account the number of tokens, hidden dimension, and numeric precision. This allows experimentation with a wide range of formats including FP4 and 1-bit signed vectors.
Figure 10: Retrieval performance (Flickr30k zero-shot image retrieval Recall@1) versus image representation storage size (in kilo-bytes, log scaled). We compare the full 576-token ViT representation and the 64-token compressed representation under BF16, FP8-E4M3, and FP4-E2M1 formats. EDJE ’s compressed tokens retain strong performance even under aggressive quantization, while substantially reducing storage.
Figure˜10 summarizes the trade-off between retrieval performance and representation storage. The plot shows that the token-compression is remarkably robust to aggressive quantization: moving from BF16 to FP8 and even FP4 leads to only minor changes in R@1, while substantially reducing storage. Moreover, combining FP8 quantization with 64-token compression cuts the per-image storage in half beyond the already compact 49 kB, with negligible loss in retrieval quality, pushing the storage-performance tradeoff even further.
Appendix I On Extending EDJE to Video Retrieval
Although video retrieval lies outside the scope of this work, EDJE naturally admits several extensions to the video domain. A video may be represented as a sequence of frames $f_{1},\dots,f_{T}]$, each processed independently by the vision encoder to produce ${\bm{X}}_{t}=g_{\phi}(f_{t})\in\mathbb{R}^{n\times d}$. Concatenating per-frame features yields a high-dimensional tensor
${\bm{\mathsfit{X}}}=[{\bm{X}}_{1},\dots,{\bm{X}}_{T}]\in\mathbb{R}^{T\times n\times d},$
corresponding to $T\times n$ visual tokens. The central challenge is therefore to compress hundreds of frames efficiently while preserving temporal structure and producing a compact representation suitable for re-ranking.
Below, we outline two concrete strategies for adapting EDJE to video-text retrieval.
  1. 1.
    EDJE’s adapter as a temporal compressor. A direct extension is to repurpose the EDJE token-compression adapter to operate across time. Each frame is first mapped to a single visual token using any image embedding model, yielding a temporal sequence of $T$ tokens. A temporal adapter can then be trained to compress these tokens into a small set of temporally aggregated tokens. This design captures long-range semantics across many frames and improves upon simple temporal pooling strategies such as averaging frame embeddings [Rasheed et al., 2022, Maaz et al., 2023]. Temporal positional encodings can be incorporated by adding a learned position vector to each frame token before passing it to the adapter.
  2. 2.
    Spatial compression followed by temporal compression. An improvement over the aforementioned idea is to enhance single frame representations by replacing naive embeddings by compressing frames via a pretrained version of EDJE’s token-compression adapter, applied to each frame to better compress information spatially. Then, a similar token-compression adapter is learned to temporally aggregate all of the compressed tokens from each frame in an efficient manner (e,g $576\rightarrow 64$ per frame).
Appendix J Impact of Text-Encoder Choice on Retrieval
The joint encoder in EDJE uses an extremely lightweight MiniLM-L12-H384 text model for joint encoding. A natural question is whether replacing MiniLM with a larger, more expressive language encoder such as BERT-Base could further improve retrieval performance. Prior work in vision language modeling suggests that stronger language encoders sometimes offer marginal gains in caption understanding, but they also introduce greater computational cost and may provide diminishing returns in retrieval settings where the visual features dominate [Devlin et al., 2019, Radford et al., 2021, Jia et al., 2021, Li et al., 2022].
To study this trade-off, we replace MiniLM with BERT-Base (uncased) while keeping all other components fixed. In particular, we use the same SigLIP2 Base visual encoder, the same token-compression module, and the same training protocol used in our main experiments. The joint encoder therefore differs only in the text backbone. We evaluate this configuration on Flickr30k using the standard zero-shot retrieval protocol.
Table 11: Flickr30k zero-shot retrieval results when substituting MiniLM with BERT-Base as the text encoder in the joint module. The visual encoder is SigLIP2-Base in all configurations. EDJE (MiniLM) denotes the baseline from the main paper.
Text Encoder Text-to-Image Image-to-Text
R@1 R@5 R@10 R@1 R@5 R@10
MiniLM with Siglip Base 84.3 96.6 97.9 94.3 99.9 99.9
BERT-Base with Siglip Base 84.52 96.58 97.9 93.9 99.9 99.9
As shown in Table 11, substituting MiniLM with BERT-Base produces mixed but notable effects. BERT-Base yields slight improvements in the text-to-image setting, particularly on R@1 and R@10 (e.g., R@1 increases from $84.3$ to $84.52$, and R@10 from $97.9$ to $97.98$). In contrast, MiniLM remains marginally stronger for image-to-text retrieval (e.g., R@1 improves from $93.9$ with BERT to $94.3$ with MiniLM), while both models tie at higher recall levels.
Although these results do not indicate that larger language encoders provide substantial gains in our setup, it is important to note that our model is trained on only 12M images, considered small-scale relative to typical vision-language pretraining regimes. Nevertheless, MiniLM, despite its compact size, does not appear to bottleneck the token-compressed retrieval pipeline.
Overall, this experiment shows that BERT-Base can slightly improve specific aspects of retrieval, but lightweight, well-distilled text encoders like MiniLM remain highly competitive and more efficient for EDJE’s joint-encoder design.